# RIS-LAD: A Benchmark and Model for Referring Low-Altitude Drone Image Segmentation

### v1

@@ V Adaptive Reasoning Fusion Module
[…]
Method Publication P@0.5 P@0.6 P@0.7 P@0.8 P@0.9 oIoU mIoU Val Test Val Test Val Test Val Test Val Test Val Test Val Test LAVT[38] CVPR 2022 33.95 31.67 31.16 27.04 26.29 23.02 19.41 17.55 9.89 8.58 44.03 41.97 32.25 30.14 ASDA[21] MM 2024 35.74 33.19 29.44 26.85 23.14 20.16 13.54 12.78 4.01 3.94 38.70 37.53 35.46 33.33 VATEX[20] WACV 2025 19.27 16.07 14.76 11.87 9.74 8.47 6.02 4.67 2.08 1.63 24.83 24.27 20.32 18.53 LGCE[26] IEEE TGRS 2024 28.22 26.38 24.07 21.82 19.99 17.55 15.11 13.32 7.88 7.06 41.72 40.75 27.68 26.17 FIANet[17] IEEE TGRS 2024 42.91 40.21 37.61 35.22 32.59 29.93 26.43 23.71 14.83 13.21 45.24 43.39 39.61 37.44 RMSIN[25] CVPR 2024 45.85 43.36 40.97 38.11 35.24 32.36 28.01 25.08 16.33 13.75 50.17 48.82 42.08 39.60 RSRefSeg[14] IGARSS 2025 47.35 44.73 44.48 40.21 39.76 36.30 35.17 30.91 26.58 21.46 50.04 47.71 43.92 41.16 CADFormer[41] JSTARS 2025 45.34 42.20 38.90 36.99 33.60 31.92 26.86 23.60 15.40 12.63 47.37 46.47 41.36 39.32 SAARN(ours) 47.06 45.02 43.12 39.34 38.25 33.37 31.81 26.24 19.27 15.31 51.54 49.60 44.30 41.67
@@
$[w_{l},w_{d},w_{c}]=\text{Softmax}(\text{SRG}(X^{fp}))$ (9)
[…]
@@
[…]
@@
VI Experiments
VI-A Implementation Details
We utilize the pre-trained Swin Transformer [36] as the visual backbone to extract multi-scale image features. For linguistic encoding, we employ the BERT-base model [37] to obtain contextual embeddings of referring expressions. The model is trained on the proposed RIS-LAD dataset, specifically curated to evaluate performance in realistic LAD environments. The model is trained for 50 epochs using the AdamW optimizer, with an initial learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. A polynomial decay schedule is employed to progressively reduce the learning rate. All experiments are conducted on four NVIDIA RTX 3080 GPUs, using a batch size of 2 per GPU.
For metrics, we report Precision@X (P@0.5–0.9), Overall Intersection-over-Union (oIoU), and Mean Intersection-over-Union (mIoU). As in previous RRSIS studies, oIoU and mIoU serve as the primary metrics for evaluating segmentation performance.
VI-B Comparison with SOTA Methods
As shown in Table II, the proposed SAARN achieves the best performance across both primary metrics: oIoU and mIoU. Although it is based on RMSIN, our approach improves oIoU from 50.17 to 51.54 and mIoU from 42.08 to 44.30 on the validation set. It further achieves a significant mIoU improvement of 2.07 on the test set. Notably, SAARN achieves substantial gains under stricter localization thresholds, including +18.0% in P@0.9 and +13.6% in P@0.8. These results demonstrate the effectiveness of the proposed method in improving semantic grounding and object discrimination in complex LAD scenarios.
Conventional RIS methods, such as ASDA and VATEX, face significant performance degradation under the LAD setting, mainly due to the small size of objects and the complex background. In particular, VATEX shows relatively limited performance, partially because it relies on a frozen CLIP image encoder for visual enhancement. This leads to negative transfer because of the difference in domain between LAD imagery and the CLIP pre-training distribution. This observation underscores the fundamental domain gap between LAD-RIS and conventional RIS.
RSRefSeg directly integrates CLIP and SAM, achieving strong results on Precision@0.5–0.9. However, qualitative results in Fig. 6 reveal that RSRefSeg tends to predict overly large masks that cover both target objects and irrelevant objects. Consequently, the high precision scores may not accurately reflect the model’s true localization ability, potentially providing a misleading indication of performance in the presence of excessive mask coverage. Compared to FIANet, which also employs linguistic feature decoupling and multi-scale integration, SAARN achieves significantly better performance, particularly on oIoU and mIoU metrics. We attribute this to the high density of same-category object distributions in LAD imagery, which exacerbates ambiguity during scale-level reasoning.
CDLE ARFM oIoU mIoU Val Test Val Test 49.77 48.32 42.08 39.60 51.31 49.70 43.97 40.68 49.82 49.28 43.31 41.02 51.54 49.60 44.30 41.67
@@ VI-B Comparison with SOTA Methods
l c d oIoU mIoU Val Test Val Test 49.82 49.28 43.31 41.02 50.64 49.44 43.52 41.43 49.90 49.30 43.47 41.17 51.06 49.52 43.98 41.61 51.54 49.60 44.30 41.67
@@ VI-C Ablation Study
We conduct ablation studies on the two core modules of SAARN, CDLE and ARFM, to validate their effectiveness. As shown in Table III, adding CDLE significantly improves mIoU. By reinforcing category semantics early in the process, CDLE helps the model focus on the correct object categories, resulting in better instance-level segmentation precision. Moreover, integrating ARFM significantly improves oIoU, increasing it from 49.82 to 51.54 on the validation set. This improvement stems from ARFM’s ability to distinguish densely packed objects across scales, enabling more accurate localization and reducing false activations.
VI-C1 Linguistic components ablation
@@
As shown in Table IV, we further conduct an analysis of the linguistic components used in the ARFM module. Specifically, we evaluate the contributions of the global, class-level and descriptive linguistic feature $l$, $c$ and $d$. Adding $l$ increases the oIoU of the validation set from 49.82 to 50.64, indicating that $l$ provides a general semantic bias that facilitates the alignment of basic characteristics across scales. Interestingly, when $d$ is further introduced with $l$, performance drops slightly. The decrease suggests that, without class-level guidance, the model becomes more susceptible to semantic interference from fine-grained descriptive cues, which may disrupt category-consistent representation learning, especially at coarser resolutions. In contrast, introducing $c$ significantly improves performance on both oIoU and mIoU, with the validation oIoU increasing by 0.42. This highlights the critical role of $c$ in semantic alignment, as it provides strong category-specific constraints. Finally, combining all three features achieves the best overall performance. These results confirm that $c$ provides essential semantic structure, while $d$ contributes complementary fine-grained cues, leading to more accurate reasoning process.
VI-D Visualization
@@
We perform a qualitative comparison with two of the SOTA RRSIS methods to provide a clearer understanding of our model’s effectiveness. As shown in Fig. 6, in typical LAD scenarios featuring tiny and densely packed objects, our model effectively identifies the correct object category. It also accurately distinguishes the target from nearby same-category instances, even under complex backgrounds and varying drone viewpoints. By contrast, selected RRSIS methods suffer from noticeable drift, including misclassification and confusion among densely clustered same-category objects.
VII Conclusion
@@
In this work, we introduced RIS-LAD, the first fine-grained dataset designed for RLADIS. Through detailed analysis, we identified two core challenges specific to this setting: category drift and object drift, both of which significantly degrade the performance of existing RES and RRSIS models. To address these issues, we propose a Semantic-Aware Adaptive Reasoning Network (SAARN). SAARN disentangles linguistic features and injects them at different stages of the network. The CDLE module strengthens early-stage category alignment to reduce semantic ambiguity, while the ARFM performs adaptive multi-scale reasoning to suppress object drift in densely distributed scenes. The experimental results demonstrate the effectiveness of our approach, showing improvements over the SOTA RIS and RRSIS methods. Our work provides a strong foundation for multi-modal understanding in complex drone-based tasks.

### v2

@@ IV-C Adaptive Reasoning Fusion Module
where $W^{q}_{s}$, $W^{k}_{s}$, and $W^{v}_{s}$ are learnable projection matrices for the query, key, and value associated with the linguistic branch $s$, and $\text{MHA}(\cdot)$ denotes the standard multi-head attention [45]. The resulting responses $\alpha_{s}$ encode the cross-modal correlation between $X^{fp}$ and each linguistic feature.
Method Publication P@0.5 P@0.6 P@0.7 P@0.8 P@0.9 oIoU mIoU
Val Test Val Test Val Test Val Test Val Test Val Test Val Test
LAVT CVPR 2022 33.95 31.67 31.16 27.04 26.29 23.02 19.41 17.55 9.89 8.58 44.03 41.97 32.25 30.14
ASDA MM 2024 35.74 33.19 29.44 26.85 23.14 20.16 13.54 12.78 4.01 3.94 38.70 37.53 35.46 33.33
VATEX WACV 2025 19.27 16.07 14.76 11.87 9.74 8.47 6.02 4.67 2.08 1.63 24.83 24.27 20.32 18.53
LGCE IEEE TGRS 2024 28.22 26.38 24.07 21.82 19.99 17.55 15.11 13.32 7.88 7.06 41.72 40.75 27.68 26.17
FIANet IEEE TGRS 2024 42.91 40.21 37.61 35.22 32.59 29.93 26.43 23.71 14.83 13.21 45.24 43.39 39.61 37.44
RMSIN CVPR 2024 45.85 43.36 40.97 38.11 35.24 32.36 28.01 25.08 16.33 13.75 50.17 48.82 42.08 39.60
RSRefSeg IGARSS 2025 46.35 44.73 43.48 40.21 39.76 36.30 35.17 30.91 26.58 21.46 50.04 47.71 43.42 41.16
CADFormer JSTARS 2025 45.34 42.20 38.90 36.99 33.60 31.92 26.86 23.60 15.40 12.63 47.37 46.47 41.36 39.32
SAARN(ours) 47.06 45.02 43.12 39.34 38.25 33.37 31.81 26.24 19.27 15.31 51.54 49.60 44.30 41.67
@@
$w_{l},w_{d},w_{c}]=\text{Softmax}(\text{SRG}(X^{fp}))$ (9)
$\text{SRG}(\cdot)$ consists of a global average pooling layer followed by two convolutional layers with ReLU activation, which produce three normalized fusion weights. The weighted attention is computed as,
@@
where $\text{Fuse}(\cdot)$ denotes the channel fusion layer, which consists of the channel concatenation followed by convolutional projection for compression and alignment.
@@
V Experiments
V-A Implementation Details
The model is trained on the proposed RIS-LAD dataset for 50 epochs using the AdamW [46] optimizer, with an initial learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. A polynomial decay schedule is employed to progressively reduce the learning rate. All experiments were performed on four NVIDIA RTX 3080 GPUs in a batch size of 8.
For metrics, we report Precision@0.5–0.9 (P@X), Overall Intersection-over-Union (oIoU), and Mean Intersection-over-Union (mIoU). While P@X highlights accurate predictions, it may overestimate performance on small objects due to their ease of enclosure. Thus, oIoU and mIoU are used as primary metrics to evaluate segmentation performance.
V-B Comparison with SOTA Methods
As shown in Table II, the proposed SAARN achieves the best performance across both primary metrics: oIoU and mIoU. Compared to RMSIN, the state-of-the-art among previous methods [36, 47] that does not incorporate pre-trained large vision models (LVMs), our approach increases oIoU from 50.17 to 51.54 on the validation set and further achieves a significant mIoU improvement of 2.07 on the test set. Notably, SAARN achieves substantial gains under stricter localization thresholds, including +18.0% in P@0.9 and +13.6% in P@0.8.
By incorporating CLIP and SAM, RSRefSeg achieves competitive performance on Precision@0.8–0.9. However, it still underperforms compared to SAARN on the core segmentation metrics. Visualization analysis reveals that the inconsistent performance of RSRefSeg due to its tendency to produce over-extended masks, as shown in Fig. 6. Although these masks may cover the referred object, their boundaries exhibit severe region overgeneralization, leading to lower oIoU and mIoU. In contrast, SAARN generates more boundary-accurate masks, resulting in stronger segmentation performance.
CDLE ARFM oIoU mIoU
Val Test Val Test
49.77 48.32 42.08 39.60
51.31 49.70 43.97 40.68
49.82 49.28 43.31 41.02
51.54 49.60 44.30 41.67
@@ V-B Comparison with SOTA Methods
l c d oIoU mIoU
Val Test Val Test
49.82 49.28 43.31 41.02
50.64 49.44 43.52 41.43
49.90 49.30 43.47 41.17
51.06 49.52 43.98 41.61
51.54 49.60 44.30 41.67
@@
Compared to FIANet, which also utilizes linguistic decoupling and multi-scale fusion, SAARN performs notably better, owing to its superior instance-level perception in dense distributions of same-class objects.
Conventional RIS methods, such as ASDA and VATEX, face significant performance degradation under the LAD setting, mainly due to the small size of objects and complex background. In particular, VATEX shows relatively limited performance, partially because it relies on a frozen CLIP image encoder for visual enhancement. This leads to negative transfer because of the difference in domain between LAD imagery and the CLIP pre-training distribution. This observation underscores the fundamental domain gap between RLADIS and conventional RIS.
@@ V-C Ablation Study
We conduct ablation studies on the two core modules of SAARN, CDLE and ARFM, to validate their effectiveness. As shown in Table III, adding CDLE significantly improves mIoU. By reinforcing category semantics early in the process, CDLE helps the model focus on the correct object categories, resulting in better instance-level segmentation precision. In addition, integrating ARFM significantly improves oIoU, increasing it from 49.82 to 51.54 on the validation set. This improvement results from ARFM’s ability to distinguish densely distributed objects across scales, facilitating more precise localization and significantly reducing false activations.
V-C1 Linguistic Components Ablation
@@
As shown in Table IV, we further conduct an analysis of the linguistic components used in the ARFM. Specifically, we evaluate the contributions of the global, class-level and descriptive linguistic feature $l$, $c$ and $d$. Adding $l$ improves the oIoU on validation set from 49.82 to 50.64, indicating its role in providing global semantic guidance to enhance alignment across multiple scales. Interestingly, performance drops slightly when $d$ is introduced alongside $l$. The decrease suggests that, without class-level guidance, the model becomes more susceptible to semantic interference from fine-grained descriptive cues, which may disrupt category-consistent representation learning, especially at coarser resolutions. In contrast, introducing $c$ significantly improves performance on both oIoU and mIoU. This highlights the critical role of $c$ in semantic alignment, as it provides strong category-specific constraints.
Finally, combining all three features achieves the best overall performance. This indicates that each linguistic component injected in ARFM plays a distinct but complementary role in the reasoning process, collectively enhancing the model’s ability to accurately localize the referred object.
V-D Visualization and Quantitative Results
@@
We present a qualitative comparison with two of the SOTA RRSIS methods to better illustrate the effectiveness of our proposed model. As shown in Fig. 6, under typical LAD scenarios, our model accurately identifies the correct object category and effectively distinguishes the object from nearby same-category instances, even in complex backgrounds and under varying drone viewpoints. In contrast, the compared RRSIS methods exhibit noticeable drift, often leading to misclassification and confusion among closely clustered objects of the same class. A detailed discussion of model limitations and failure cases is provided in the appendix.
VI Conclusion
@@
In this work, we propose RIS-LAD, the first fine-grained dataset designed for RLADIS. Through detailed analysis, we identified two core challenges specific to this setting: category drift and object drift, both of which significantly degrade the performance of existing RES and RRSIS models. To address these issues, we propose a semantic-aware adaptive reasoning network which comprises two key modules: CDLE and ARFM. CDLE enhances early-stage category alignment, while ARFM performs adaptive multi-scale reasoning to mitigate object drift. The experiments demonstrate the strong performance of our framework and underscore the challenges posed by RIS-LAD, setting a new benchmark for LAD scenarios.